OpenStudy (ivyjane):

in Angle ABC, Angle A = 42. The measure of Angle B is Five times the meausre of angle C. Find angle B and Angle C

OpenStudy (anonymous):

\[180 - 42 = 138 = B + C\] \[B = 5C\] \[138 = (5C) + C = 6C\] \[\frac{ 138 }{ 6 } = C = 23\] \[B = 5(23) = 115\] B is therefore 115 and C is 23.

OpenStudy (ivyjane):

thanks!!!!!